# Humanoid Whole-Body Badminton via Multi-Stage Reinforcement Learning

### v1

@@ Abstract
Humanoid robots have demonstrated strong capability for interacting with deterministic scenes across locomotion, manipulation, and more challenging loco-manipulation tasks. Yet the real world is dynamic, quasi-static interactions are insufficient to cope with the various environmental conditions. As a step toward more dynamic interaction scenario, we present a reinforcement-learning–based training pipeline that produces a unified whole-body controller for humanoid badminton, enabling coordinated lower-body footwork and upper-body striking without any motion priors or expert demonstrations. Training follows a three-stage curriculum—first footwork acquisition, then precision-guided racket swing generation, and finally task-focused refinement—yielding motions in which both legs and arms serve the hitting objective. For deployment, we incorporate an Extended Kalman Filter (EKF) to estimate and predict shuttlecock trajectories for target striking. We also introduce a prediction-free variant that dispenses with EKF and explicit trajectory prediction. To validate the framework, we conduct five sets of experiment in both simulation and the real world. In simulation, two robots sustain a rally of 21 consecutive hits. Moreover, the prediction-free variant achieves successful hits with comparable performance relative to the target-known policy. In real-world tests, both the prediction and controller module exhibit high accuracy, and on-court hitting achieves an outgoing shuttle speed up to 10 m/s with a mean return landing distance of 3.5 m. These experiment results show that our humanoid robot can deliver highly dynamic while precise goal striking in badminton, and can be adapted to more dynamism critical domains.
@@ I Introduction
Humanoid platforms have been proposed as general-purpose embodied agents for human-compatible skills [1, 2]. Despite rapid progress in locomotion and motion imitation, agile, contact-rich interactions with fast-moving objects under tight reaction windows remain underexplored. Badminton is a particularly ideal testbed: returns require sub-second perception–action loops, precisely timed and oriented racket–shuttle contacts within a large 3D interception volume, and whole-body coordination that blends rapid arm swings with stable and agile leg motions.
@@
Compared with several recent works on robot table tennis [3, 4, 5], badminton aggravates several difficulties. Aerodynamics are highly uncertain: strong drag and a non-parabolic “flip” regime make the pre-turning trajectory hard to model; as a result, the seemingly longer flight still yields less usable decision time. Although the racket face is larger, valid hits occur only in brief windows when position, velocity, and orientation align; swing amplitudes are also much larger, injecting angular momentum that challenges whole-body balance. Critically, footwork and striking must co-evolve: lower-body motion does not merely reposition the base but co-determines hitting accuracy. Our system reaches swing speeds around 5.3 m/s and produces outgoing shuttle speeds up to 10 m/s. The serve-to-hit window is 0.8 to 1.0 s; the first 0.36 s is consumed by trajectory prediction, leaving less than 0.6 s for the controller to react.
@@
Humanoids are a natural fit for badminton. Lateral mobility and reach patterns match human-centered rules and court geometry, and established human strategies (e.g., split steps, approach phases, recovery) transfer more directly compared to the non-humanoid platforms. Demonstrating badminton on a humanoid, therefore, serves as a stepping stone toward broader, high-dynamics, and interactive behaviors in human environments.
@@
Related efforts highlight both promise and gaps. [5] demonstrated humanoid table tennis via a hierarchical planner coupled to a learned whole-body controller and achieved long rallies. Our setting differs in key ways: (i) we do not use reference motions to shape striking; instead, the policy discovers energy-efficient swings purely via reinforcement learning, simplifying implementation, and avoiding excessive style constraints; (ii) The use of “virtual hit plane” reduces striking to a largely 2D problem, whereas badminton demands orientation-aware contacts throughout a 3D free space; and (iii) badminton’s large-amplitude strokes induce significant whole-body disturbance, requiring balance recovery that table-tennis-scale motions seldom provoke. Closer to badminton, recent work on legged mobile manipulation learns coordinated skills with onboard vision [6]; however, humanoid hardware deployment remained elusive, and their simulated policies did not exhibit badminton-style footwork, suggesting missing training signals for lower-body coordination. These gaps motivate a training pipeline that explicitly fosters footwork–strike synergy on humanoids and scales to real hardware.
@@
This paper presents, to our knowledge, the first real-world humanoid robot that plays badminton. We train a unified whole-body controller with a multi-stage reinforcement learning curriculum: (1) footwork acquisition toward target regions with feet-swing time and height regularization; (2)precision-guided arm swing generation with timing and scheduled tightening of position and orientation accuracy; and (3) final refinement that removes most locomotion-shaping regularizers to avoid gradient interference and maximize hitting performance. Across all stages, we adopt the shuttlecock simulation of [6] to generate flight trajectories for training. At each rollout, the policy receives a compact target tuple—hit time, intercept position, and racket orientation—drawn from a pre-generated corpus of about 20k trajectories; intercepts are selected randomly within a 1.5–1.6 m height band. Exponential rewards with scheduled tightness transition behavior from coarse to precise.
@@
The policy observes target hit information as well as proprioception and inferred actions for 21 joints on a 1.28 m custom humanoid. In deployment, the hitting target is obtained online from an EKF that predicts the shuttlecock trajectory. We leverage a Low-level PD controller that runs at 500 Hz, with policy inference at 50 Hz to derive the desired joint position. Training occurs fully in simulation with moderate domain randomization, and the final control policy is deployed in a zero-shot manner without system identification or manual parameter adjustment. We conduct the real-world experiments in a motion capture (Mocap) arena with a serving machine.
@@
Beyond the default pipeline, we also study a prediction-free variant that removes the explicit shuttle-trajectory predictor. During training, the actor only receives instantaneous shuttle pose plus five history pose frames to infer hit timing and pose; the critic retains privileged targets (asymmetric actor–critic) to better learn the value function [7]. This end-to-end control strategy aligns with human intuition (rolling anticipation), simplifies deployment by avoiding predictor or planner retuning, and naturally supports aerodynamic randomization. We report simulation-base comparison results in this work while leave real-robot validation to the future work.
@@
Both simulation and real world results indicate that our humanoid can execute large-amplitude swings while maintaining balance, react within sub-second windows, and intercept fast incoming shuttles fully autonomously without the aid of teleportation or motion demonstrations. While our current tests focus on machine-served shuttles in the Mocap arena, observed behaviors from the training recipe suggest a feasible path to higher-ceiling venues and, ultimately, rallies with robotic or human opponents. The key technical contributions are summarized as follows:
@@
• First real-world humanoid badminton: an autonomous, unified whole-body controller that returns machine-served shuttles on a 21-DoF humanoid, achieving about 5.3 m/s swing speeds and producing outgoing shuttle speeds up to 10 m/s under sub-second reaction.
@@
• A prediction-free variant: a more end-to-end policy that infers timing and hitting target implicitly from short-horizon shuttle observations, improving robustness to aerodynamic variability and simplifying deployment by removing EKF-based trajectory predictor.
@@ II Related Work
II-A Mobile Whole-Body Control
Combining manipulation skills with flexible mobile ability is essential for a wide range of robotic applications. Typically, the base movement and end-effector tracking problems can be solved separately to reduce task complexity [8, 9]. The work in [9] proposes an MPC controller to track the desired end-effector position of an arm mounted on the quadruped with a standalone RL-based controller for locomotion. To catch objects in flight, [10] assigns the base movement, arm swing, and hand catching tasks to three policy modules. Similarly, [8] claims that decoupling upper-body and lower-body control into separate agents can achieve a more precise end-effector tacking performance when the robot is moving.
Despite these promising advantages, there are still many efforts made to unify the lower and upper body control [11, 12, 13]. Specifically, [11, 12], [14] leverage a multi-critic architecture to ease the optimization of the coordinated behavior. Moreover, [13] introduces a physical feasibility reward design to guide the unified policy learning. As for recent works on dynamic whole-body tasks such as racket sports and tossing, performance hinges on coordinated whole-body impulse generation and timing, with the legs contributing to power transmission rather than merely locomotion. Unified whole-body control has proved effective across these domains [15], [6], [16].
@@ II-B Robotic Racket Sports
Devising control strategies for robotic Racket Sports has been a fundamental research benchmark that helps inspect progress on the whole-body control algorithms [5, 3, 17, 6]. The intensive coupling of environmental perception, target reaching, and precise end-effector control brings a critical challenge to the system design. To achieve an amateur human-level table tennis competition, [3] introduces a hierarchical and modular policy architecture to handle the unseen scenarios when playing with human. Despite the impressive result in [3], [5] shows that a reinforcement learning-based whole-body controller with a simple model-based planner for ball trajectory prediction and racket target planning are competent for a humanoid robot to conduct consecutive shots with a human opponent. More recently, [6] introduces a perception-informed whole-body visuomotor policy that achieves effective shuttlecock tracking and striking using a quadrupedal platform for the badminton task. However, we assume that the humanoid robot platform shows a more consistent motion pattern with humans for the badminton task. In this work, we focus on the coordinated badminton skill learning on a self-developed humanoid platform and conduct a systematic evaluation of the controller’s performance.
@@ III System Overview
Platform. Our robot is a self-developed general-purpose humanoid Phybot C1 (1.28 m, 30 kg, 21 DoF), see Fig. 2 (b). DoF distribution: each leg has 3 hip joints, 1 knee, 2 ankle; the torso has 1 waist; each arm has 3 shoulder joints and 1 elbow. All joints use cycloidal-reduction actuators. Lower-body peak torques span 96–244 Nm; most arm joints are 96 Nm except shoulder-yaw (12 Nm). Motors reach maximum 15 rad/s. A standard 100 g full size badminton racket is rigidly clamped to the distal forearm by a 3D-printed mount, with the racket perpendicular to the forearm axis. Onboard sensing comprises an IMU and absolute encoders; we do not employ force or tactile sensors.
@@
External devices and power. Serves are produced by a manually aimed ball machine with adjustable position and launching angle, up to 14 m/s shuttle speed. The robot is powered by a 72 V battery pack. Further hardware specifications (joint limits, torque constants, detailed parameters) are provided in the appendix.
@@ IV-A1 Multi-Stage Reinforcement Learning
Observation space. The actor’s basic observation is 87-D, concatenating proprioception information and external sensing from Mocap. On top of this 87-D vector, we stack short/long history of all joint positions, velocities, and actions (5 frames + 20 frames) [18], adding up to 1,557 features to mitigate partial observability (e.g., base-velocity noise). The critic observation has 98 dimensions in total, which include the privileged information used to enhance actor version observations, such as noise-free base and joint states as well as racket speed which may be difficult to acquire in the real world. Others are classified as preemptive knowledge used to make MDP well-posed for accurate value learning. All quantities are expressed in the world frame. Detailed information are listed in Fig. 2(a), (c).
@@
Episodes Settings. Each episode contains six swing targets, following the configuration used by [6]. This design enables follow-through behavior learning but makes the value function depend on the number of hits remaining. We therefore adopt an asymmetric actor–critic [19] the critic receives additional preemptive information as mentioned above to stabilize value estimation, while the actor only uses deployable observations. Concretely, these are two subsequent hitting times, next target position and orientation, and the number of remaining targets.
@@
  • S3 — Task-focused refinement: remove target approach rewards (which is the main task reward in S1) and most gait-shaping regularization (e.g. foot distance, step/contact symmetry, step height) to avoid gradient interference with the hitting objective; keep safety, energy, hardward limit constraint terms. Domain randomization and observation noise are enabled here to consolidate robustness.
We conduct a simple ablation study. Removing S1 and starting from S2 causes training to diverge directly. Using S1–S2 yields a hardware-deployable policy but there’s still room for improvement. S3 breaks this plateau and improves performance and robustness.
@@ IV-A2 Multi-Stage Reward Design
We devise a modular reward function that consists of a locomotion-style term $r_{\mathrm{loco}}$ and an arm hitting term $r_{\mathrm{hit}}$ to guide the emergence of coordinated badminton skills. The locomotion reward follows [22] and leverages feet swing-time to encode step frequency implicitly. While the hit reward references [6] but making some modifications. We group the remaining terms into global regularization (action rate, joint limits, collision, energy etc.). Exact weights are reported in the appendix.
@@
S2 — Precision-guided swing generation. On top of S1, we lower the weight of regional tracking and introduce a hit-instant reward with large weight (six activations per episode). This reward comes with two terms, hitting precision and racket swinging speed. While [6] splits hitting precision into position and orientation with different scale, we find it to be more accurate to entangle two terms together, coupling position and orientation accuracy [15]. Then we introduce another racket speed term to encourage swing. Define the target racket normal $n^{*}$ as the negative $z$-axis of target hitting orientation $q^{*}_{\mathrm{ee}}$. Let $v_{\mathrm{ee}}$ be the end-effector linear velocity; the effective speed component in the direction of target is $v_{\mathrm{ee}}\!\cdot n^{*}$ (clamped at $\geq 0$). Let $e_{ee\_pos}=\|p_{\mathrm{ee}}-p^{*}_{\mathrm{ee}}\|$ and $e_{ee\_ori}=\mathrm{ang\_err_{z}}(q_{\mathrm{ee}},q^{*}_{\mathrm{ee}})$. The hit reward, active only at $t=t^{*}$, is
@@
which encourages the policy to adopt a more human-like hitting posture—performing a backswing and then swinging forward along the reverse of that backswing—to produce a more complete kinetic chain. And reward a default holding pose when no shuttle is launched,
@@
S3 — Task-focused refinement. Starting from the S2 checkpoint, we remove most gait-shaping terms in $r_{\mathrm{loco}}$—foot distance, contact and step symmetry, optimal footholds, step height—to avoid gradient conflict with the hitting objective. We retain gloable regularization terms and $r_{\mathrm{hit}}$ rewards, and enable domain randomization plus observation noise to consolidate robustness. With these enabled—thus making the environment deliberately harder—the policy does not regress; instead, we observe consistent improvements across metrics. The primary hitting reward $r_{swing}$ increases by $3\text{--}5\%$, and we see lower action rates, reduced joint velocity and acceleration penalties, lower energy consumption, cleaner foot-contact force profiles, and reduced joint-torque usage; notably, the energy and torque costs decrease by $\approx 20\%$ empirically.

### v2

@@ Abstract
Humanoid robots have demonstrated strong capabilities for interacting with static scenes across locomotion, manipulation, and more challenging loco-manipulation tasks. Yet the real world is dynamic, and quasi-static interactions are insufficient to cope with diverse environmental conditions. As a step toward more dynamic interaction scenarios, we present a reinforcement-learning–based training pipeline that produces a unified whole-body controller for humanoid badminton, enabling coordinated lower-body footwork and upper-body striking without motion priors or expert demonstrations. Training follows a three-stage curriculum—first footwork acquisition, then precision-guided racket swing generation, and finally task-focused refinement—yielding motions in which both legs and arms serve the hitting objective. For deployment, we incorporate an Extended Kalman Filter (EKF) to estimate and predict shuttlecock trajectories for target striking. We also introduce a prediction-free variant that dispenses with EKF and explicit trajectory prediction. To validate the framework, we conduct five sets of experiments in both simulation and the real world. In simulation, two robots sustain a rally of 21 consecutive hits. Moreover, the prediction-free variant achieves successful hits with comparable performance relative to the target-known policy. In real-world tests, both prediction and controller modules exhibit high accuracy, and on-court hitting achieves an outgoing shuttle speed up to 19.1 m/s with a mean return landing distance of 4 m. These experimental results show that our proposed training scheme can deliver highly dynamic while precise goal striking in badminton, and can be adapted to more dynamics-critical domains.
@@ I Introduction
Humanoid platforms have been proposed as general-purpose embodied agents for human-compatible skills [1, 2, 3, 4, 5, 6, 7]. Despite rapid progress in locomotion and motion imitation, agile, contact-rich interactions with fast-moving objects under tight reaction windows remain underexplored. Badminton is a particularly ideal testbed: returns require sub-second perception–action loops, precisely timed and oriented racket–shuttle contacts within a large 3D interception volume, and whole-body coordination that blends rapid arm swings with stable and agile leg motions.
@@
Compared with several recent works on robotic table tennis [8, 9, 10, 11], badminton aggravates several difficulties. Aerodynamics are highly uncertain: strong drag and a special “flip” regime make the pre-turning trajectory hard to model; as a result, the seemingly longer flight still yields less usable decision time. Although the racket face is larger, valid hits occur only in brief windows when position, velocity, and orientation align; swing amplitudes are also much larger, injecting angular momentum that challenges whole-body balance. Critically, footwork and striking must co-evolve: lower-body motion does not merely reposition the base but co-determines hitting accuracy. Our system reaches swing speeds around 5.3 m/s and produces outgoing shuttle speeds up to 19.1 m/s. The serve-to-hit window is 0.8 to 1.0 s; the first 0.36 s is consumed by trajectory prediction, leaving less than 0.6 s for the controller to react.
@@
Humanoids are a natural fit for badminton. Lateral mobility and reach patterns match human-centered rules and court geometry, and established human strategies (e.g., split steps, approach phases, recovery) transfer more directly compared to non-humanoid platforms. Demonstrating badminton on a humanoid, therefore, serves as a stepping stone toward broader, highly dynamic, interactive behaviors in human environments.
@@
Related efforts highlight both promise and gaps. [10] demonstrated humanoid table tennis via a hierarchical planner coupled to a learned whole-body controller and achieved long rallies. Our setting differs in key ways: (i) we do not use reference motions to shape striking; instead, the policy discovers energy-efficient swings purely via reinforcement learning, simplifying implementation, and avoiding excessive style constraints; (ii) The use of “virtual hit plane” reduces striking to a largely 2D problem, whereas badminton demands orientation-aware contacts throughout a 3D space; (iii) badminton’s large-amplitude strokes induce significant whole-body disturbance, requiring balance recovery that table-tennis-scale motions seldom provoke; and (iv) we do not explicitly compute target base positions; instead, the policy only receives the desired interception target and must coordinate the whole-body motion. As for badminton, recent work on a legged-manipulator learns coordinated skills with onboard vision [12]; however, humanoid hardware deployment has remained elusive, and their simulated humanoid badminton policies did not exhibit badminton-style footwork, suggesting missing training signals for lower-body coordination. These gaps motivate a training pipeline that explicitly fosters footwork–strike synergy on humanoids and scales to real hardware.
@@
This paper presents, to our knowledge, the first real-world humanoid robot that plays badminton. We train a unified whole-body controller with a multi-stage reinforcement learning curriculum: (1) footwork acquisition toward target regions with feet-swing time and height regularization; (2) precision-guided arm swing generation with timing and scheduled tightening of position and orientation accuracy; and (3) final refinement that removes numerous locomotion-shaping regularizers to avoid gradient interference and maximize hitting performance. Across all stages, we adopt the shuttlecock simulation of [12] to generate flight trajectories for training. At each rollout, the policy receives a compact target tuple—hit time, intercept position, and target racket orientation—drawn from a pre-generated corpus of about 20k trajectories; intercepts are selected randomly within a 1.5–1.6 m height band, chosen based on the humanoid’s height as a natural overhead hitting zone. Exponential rewards with scheduled tightness transition behavior from coarse to precise.
@@
The policy observes target hit information as well as proprioception and inferred actions for 21 joints on a 1.28 m custom humanoid. In deployment, the hitting target is obtained online from an EKF that predicts the shuttlecock trajectory. We leverage a Low-level PD controller that runs at 500 Hz, with policy inference at 50 Hz to derive the desired joint positions. Training occurs fully in simulation with moderate domain randomization, and the final control policy is deployed in a zero-shot manner without system identification or manual parameter adjustment. We conduct real-world experiments in a motion capture (Mocap) arena with a serving machine.
@@
Beyond our EKF-based pipeline, we also explore a prediction-free variant that removes any explicit shuttle predictor. In contrast to recent humanoid table tennis systems where end-to-end RL is combined with an auxiliary learnable predictor that outputs hitting targets for a downstream controller [11], our variant does not introduce a separate prediction module. During training, the actor only receives the instantaneous shuttle pose plus five history pose frames to infer hit timing and target pose, while the critic retains privileged target to better learn the value function. This more tightly integrated end-to-end control strategy is conceptually closer to human-like rolling anticipation and may simplify deployment by avoiding additional predictor or planner tuning, but in this paper we only provide initial simulation evidence that the system can tolerate the absence of an explicit shuttle predictor. In the remainder of this work, we therefore focus on the EKF-based pipeline as our primary method, and report simulation-based comparisons to the prediction-free variant, leaving real-robot validation of this variant to future work.
@@
Both simulation and real-world results indicate that our humanoid can execute large-amplitude swings while maintaining balance, react within sub-second windows, and intercept fast incoming shuttles fully autonomously without the aid of teleportation or motion demonstrations. While our current tests focus on machine-served shuttles in the Mocap arena, observed behaviors from the training recipe suggest a feasible path to higher-ceiling venues and, ultimately, rallies with robotic or human opponents. The key technical contributions are summarized as follows:
@@
• First real-world humanoid badminton: an autonomous, unified whole-body controller that returns machine-served shuttles on a 21-DoF humanoid, achieving about 5.3 m/s swing speeds and producing outgoing shuttle speeds up to 19.1 m/s under sub-second reaction.
@@
• A prediction-free variant: a more end-to-end policy that infers timing and hitting target implicitly from short-horizon shuttle observations, improving robustness to aerodynamic variability and removing explicit trajectory predictor.
@@ II Related Work
II-A Whole-Body Loco-manipulation
Combining manipulation skills with flexible mobility is essential for a wide range of robotic applications. Typically, the base movement and end-effector tracking problems can be solved separately to reduce task complexity [13, 14, 15, 16]. The work in [13] proposes an MPC controller to track the desired end-effector position of an arm mounted on the quadruped with a standalone RL-based controller for locomotion. To catch objects in flight, [14] assigns the base movement, arm swing, and hand catching tasks to three policy modules. Similarly, [15] claims that decoupling upper-body and lower-body control into separate agents can achieve a more precise end-effector tracking performance when the robot is moving. Despite these promising advantages, there have been many efforts to unify lower and upper body control. Specifically, [17, 18, 19] leverage a multi-critic architecture to ease the optimization of coordinated behavior. Moreover, [20] introduces a physical feasibility reward design to guide the unified policy learning. As for recent works on dynamic whole-body tasks such as racket sports and tossing, performance hinges on coordinated whole-body impulse generation and timing, with the legs contributing to power transmission rather than merely locomotion. Unified whole-body loco-manipulation has proved effective across these domains [12, 21, 22].
@@ II-B Robotic Racket Sports
The development of control strategies for robotic racket sports has become a key benchmark for assessing progress in whole-body loco-manipulation [10, 11, 12, 16, 23, 24]. The tight coupling between perception, target interception, and precise end-effector control makes these tasks particularly challenging. To achieve human–level table tennis, [8] proposes a highly modular and hierarchical policy architecture, combining learned components with carefully engineered modules. However, the resulting system is complex, requires significant manual design and tuning across modules, and relies heavily on real-world data collection. Building on this line of work, [10] and [11] both tackle humanoid table tennis. [10] demonstrates that a RL-based whole-body controller, together with a simple model-based planner for ball trajectory prediction and racket target planning, is sufficient for a humanoid robot to sustain rallies with a human opponent. Yet, this work depends on expert demonstrations to construct high-quality reference motions. Meanwhile, [11] jointly trains an additional module during RL to predict ball trajectories and propose racket targets. In both cases, however, the base motion is executed by a separate command rather than emerging from a single unified goal, so lower-body footwork and upper-body striking are not jointly optimized by the policy. Moreover, [12] introduces a perception-informed whole-body policy for badminton on a quadrupedal platform, achieving impressive shuttlecock tracking and striking performance. However, the quadruped morphology substantially simplifies balance compared to a humanoid with a narrow support polygon and a high center of mass, making the coordination of balance and large-amplitude racket swings considerably less challenging. In this work, we instead focus on coordinated badminton skill learning on a humanoid platform. We train a single whole-body controller via a multi-stage reinforcement-learning curriculum, without expert demonstrations or a separate base command, and conduct a systematic evaluation of its performance in both simulation and real-world experiments.
@@ III System Overview
Platform. The robot used in this work is a general-purpose humanoid Phybot C1 (1.28 m, 30 kg, 21 DoF), see Fig. 2(b). DoF distribution: each leg has 3 hip joints, 1 knee, 2 ankles; the torso has 1 waist; each arm has 3 shoulder joints and 1 elbow. All joints use cycloidal-reduction actuators. A standard 100 g full size badminton racket is rigidly clamped to the distal forearm by a 3D-printed mount, with the racket perpendicular to the forearm axis. Onboard sensing comprises an IMU and absolute encoders; we do not employ force or tactile sensors. Further hardware specifications are provided in the appendix.
@@
External devices and power. Serves are produced by a manually aimed ball machine with adjustable position and launching angle, up to 14 m/s shuttle speed. The robot is powered by a 72 V battery pack.
@@ IV-A1 Multi-Stage Reinforcement Learning
Observation space. The actor’s basic observation is 87-D, concatenating proprioception information and external sensing from Mocap. On top of this 87-D vector, we stack short/long history of all joint positions, velocities, and actions (5 frames + 20 frames) [25], adding another 1470 features to mitigate partial observability (e.g., base-velocity noise). The critic observation has 98 dimensions in total, which include the privileged information used to enhance actor version observations, such as noise-free base and joint states as well as racket speed which may be difficult to acquire in the real world. Others are classified as preemptive knowledge used to make MDP well-posed for accurate value learning. All quantities are expressed in the world frame. Detailed information is listed in Fig. 2(a), (c).
@@
Episode Settings. Each episode contains six swing targets, following the configuration used by [12]. This design enables follow-through behavior learning but makes the value function depend on the number of hits remaining. We therefore adopt an asymmetric actor–critic [26], the critic receives additional preemptive information as mentioned above to stabilize value estimation, while the actor only uses deployable observations. Concretely, these are two subsequent hitting times, next target position and orientation, and the number of remaining targets.
@@
  • S3 — Task-focused refinement: remove target approach reward (which is the main task reward in S1) and numerous gait-shaping regularization (e.g. foot distance, step/contact symmetry, step height) to avoid gradient interference with the hitting objective; keep safety, energy, and hardward limit constraint terms. Domain randomization and observation noise are enabled here to consolidate robustness.
We conduct a simple ablation study. Removing S1 and starting from S2 causes training to diverge directly. Skipping S2 and jumping directly from S1 to S3 makes the curriculum gap too large, and training again fails to reliably converge. Using S1–S2 yields a hardware-deployable policy but there’s still room for improvement. S3 breaks this plateau and improves performance and robustness.
@@ IV-A2 Multi-Stage Reward Design
We devise a modular reward function that consists of a locomotion-style term $r_{\mathrm{loco}}$ and an arm hitting term $r_{\mathrm{hit}}$ to guide the emergence of coordinated badminton skills. The locomotion reward follows [29] and leverages feet swing-time to encode step frequency implicitly. While the hit reward references [12] but making some modifications. We group the remaining terms into global regularization $r_{\mathrm{global}}$ (action rate, joint limits, collision, energy etc.). Exact weights are reported in the appendix.
@@
S2 — Precision-guided swing generation. On top of S1, we lower the weight of regional tracking and introduce a hit-instant reward with large weight (six activations per episode). This reward comes with two terms, hitting precision and racket swinging speed. While [12] splits hitting precision into position and orientation with different scales, we find it to be more accurate to entangle two terms together, coupling position and orientation accuracy [21]. Then we introduce another racket speed term to encourage swing. Define the target racket normal $n^{*}$ as the $z$-axis of target hitting orientation $q^{*}_{\mathrm{ee}}$. Let $v_{\mathrm{ee}}$ be the end-effector linear velocity; the effective speed component in the direction of target is $v_{\mathrm{ee}}\!\cdot n^{*}$. Let $e_{ee\_pos}=\|p_{\mathrm{ee}}-p^{*}_{\mathrm{ee}}\|$ and $e_{ee\_ori}=\mathrm{ang\_err_{z}}(q_{\mathrm{ee}},q^{*}_{\mathrm{ee}})$ (defined as the angle between the current and target racket normals; only the normal direction is considered here, rewarding the full EE orientation would be possible but is not required for the hitting task). The hit reward, active only at $t=t_{hit}$, is
@@
which encourages the policy to adopt a more human-like hitting posture—performing a backswing and then swinging forward along the reverse of that backswing—to produce a more complete kinetic chain. We also reward a default holding pose when no shuttle is launched,
@@
S3 — Task-focused refinement. Starting from the S2 checkpoint, we remove target approach reward and numerous gait-shaping terms in $r_{\mathrm{loco}}$—foot distance, contact and step symmetry, step height—to avoid gradient conflict with the hitting objective. We retain global regularization terms and $r_{\mathrm{hit}}$ rewards, and enable domain randomization plus observation noise to consolidate robustness. With these enabled—thus making the environment deliberately harder—the policy does not regress; instead, we observe consistent improvements across metrics. The primary hitting reward $r_{swing}$ increases by $3\text{--}5\%$, and we see lower action rates, reduced joint velocity and acceleration penalties, lower energy consumption, cleaner foot-contact force profiles, and reduced joint-torque usage; notably, the energy and torque costs decrease by $\approx 20\%$ empirically.